OpenStudy (macielynn):

x-y+3z=19 -2x+2y-6z=9 3x+5y+2z=3

OpenStudy (anonymous):

LINE2: (-2x+2y-6z=9)/(-2) ==> x-y+3z=(9/2) LINE1: x-y+3z=19 so it doesn't have any answer.

OpenStudy (anonymous):

I am sorry I do not know

OpenStudy (mathmate):

The system is inconsistent, and thus has no solution. Can you please post the image of the complete question so we know what you're supposed to do with the system of equations?